# Resource Allocation in Distributed Quantum Computing Interconnect Networks

### v1

@@ V Evaluation and results
Figure 9 illustrates $C_{\rm com}$, $N_{\rm part}$, and $N_{\rm rgate}$, normalised by $r_{\rm asg}$. The results clearly demonstrate that the proposed method surpasses the benchmark method in terms of inter-QPU communication cost and the number of circuit partitions. The normalised average number of remote gates is slightly higher for the proposed method compared to the benchmark, which can be attributed to the lower assignment of QFT circuits in the latter. Comparing the proposed method with and without the limitation on $K^{\rm max}$, we observe that the normalised inter-QPU communication cost is reduced by around $80\%$, while the assignment ratio decreases by only $5\%$. This highlights the trade-off between these two parameters. The network controller can intelligently explore and consider this trade-off to determine whether limiting $K^{\rm max}$ is beneficial in specific scenarios.
@@ VI Conclusion
In this work, we addressed the challenge of efficient QPU allocation and QCirc scheduling in DQC networks. We proposed a multi-objective optimisation algorithm for QPU allocation that minimises the impact of inter-QPU communication latencies while maximising the number of assigned QCircs. The proposed optimisation algorithm considers the topology of the quantum network, link latencies, the properties and limitations of the QPUs, and the inherent features of the QCircs to effectively assign resources. We evaluated the proposed algorithm using various simulations. Our Results demonstrate that the proposed method efficiently allocates computing resources, enabling faster and more accurate execution of quantum circuits. In addition to our novel QPU allocation algorithm, we present an efficient QCirc scheduling model that seamlessly complements the allocation strategy.
In this work, we primarily focused on DQC networks that leverage both quantum and classical channels to generate entanglement between QPUs. However, alternative quantum computational techniques, such as circuit knitting, has been proposed in the literature, that require only classical channels for communication between QPUs [29, 30]. One exciting future research direction is to explore QPU allocation strategies for DQC networks that either employ solely classical channels or utilise both techniques. Another potential area of investigation is the integration of QPU allocation with advanced quantum compilation techniques, such as the noise-adaptive compilation method proposed in [5].

### v2

@@ V Evaluation and results
Figures 9 (a), (b), and (c) illustrate the scaled communication cost ($C_{\rm com}^{(s)}$), scaled partition count ($N_{\rm rgate}^{(s)}$), and scaled remote gate count ($N_{\rm rgate}^{(s)}$), respectively, for two scenarios: one where the partition count for the QFT circuit is limited to one ($K^{\rm max}(\rm QFT)=1$), and another where there is no limitation on $K^{\rm max}(\rm QFT)$. The results clearly demonstrate that the proposed method surpasses the benchmark method in terms of inter-QPU communication cost and the number of circuit partitions. The scaled remote gate count is slightly higher for the proposed method compared to the benchmark, which can be attributed to the lower assignment of QFT circuits in the latter. Comparing the proposed method with and without the limitation on $K^{\rm max}$, we observe that the scaled communication cost is reduced by around $80\%$, while the assignment ratio decreases by only $5\%$. This highlights the trade-off between these two parameters. The network controller can intelligently explore and consider this trade-off to determine whether limiting $K^{\rm max}$ is beneficial in specific scenarios.
@@ VI Conclusion
In this work, we addressed the challenge of efficient QPU allocation and QCirc scheduling in DQC networks. We proposed a multi-objective optimisation algorithm for QPU allocation that minimises the impact of inter-QPU communication latencies while maximising the number of assigned QCircs. The proposed optimisation algorithm considers the topology of the quantum network, link latencies, the properties and limitations of the QPUs, and the inherent features of the QCircs to effectively assign resources. We evaluated the proposed algorithm using various simulations. Our Results demonstrate that the proposed method efficiently allocates computing resources, enabling more accurate execution of quantum circuits and improved resource utilization. In addition to our novel QPU allocation algorithm, we present an efficient QCirc scheduling model that seamlessly complements the allocation strategy.
In this work, we primarily focused on DQC networks that leverage both quantum and classical channels to generate entanglement between QPUs. However, alternative quantum computational techniques, such as circuit knitting, has been proposed in the literature, that require only classical channels for communication between QPUs [29, 30]. One exciting future research direction is to explore QPU allocation strategies for DQC networks that either employ solely classical channels or utilise both techniques. Other promising areas for future research include developing effective priority assignment policies and more efficient scheduling methods that account for the execution time of quantum circuits.